# An Introduction to Reinforcement Learning: Fundamental Concepts and Practical Applications

### v1

@@ 10 Conclusion
In this paper, we have introduced the fundamental concepts and methodologies of Reinforcement Learning (RL) in an accessible manner for beginners. We have established a foundation for understanding how RL agents learn and make decisions by providing a detailed description of the core elements of RL, such as states, actions, policies, and reward signals. The purpose of this paper is to provide an overview of various RL algorithms, including both model-free and model-based approaches, in order to illustrate the diversity of approaches within the field of RL. To facilitate deeper learning and practical application of RL, this guide aims to provide new learners with the necessary knowledge and confidence to embark on their RL journey. As part of future research, we will analyze RL in the literature in order to identify its strengths and weaknesses, providing a valuable resource for researchers. In addition, a benchmark dataset of RL algorithms will be developed from various sources, providing detailed information on RL papers and their applications.
[…]

### v2

@@ 6 Conclusion
This paper presents an introductory exploration of the fundamental concepts and methodologies of Reinforcement Learning (RL), tailored to beginners. It establishes a foundational understanding of how RL agents learn and make decisions by thoroughly examining key components such as states, actions, policies, and reward signals. By analyzing Multi-armed bandit problem, this paper introduced background of RL in an accessible and easy-to-understand way. The primary objective is to offer an overview of a wide range of RL algorithms, encompassing both model-free and model-based approaches, thereby highlighting the diversity within the field. Through this guide, we aim to equip new learners with the essential knowledge and confidence to begin their journey into RL.